an alpha particle of momentum p is bombarded on the nucleus the distance of the closest aproch is r if momentum of alpha particle is made 6 times the distance of the closest approch becomes

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

let m be mass of alpha particle. p2/2m =k/r, k is a constant, r/k=2m/p2

Thus, r is inversely proportional to (p^2), therefore if momentum of alpha particle is multiplied by 6, r is divided by 36